The second angle of a triangle measure 10 more than the first angle. The third angle measures 30 less than twice the second angle. what are the measure of the 3 angles
Answer by stanbon(75887)

The second angle of a triangle measure 10 more than the first angle. The third angle measures 30 less than twice the second angle. what are the measure of the 3 angles
----
1st angle: x
2nd angle: x+10
3rd angle: 2x-30
----
Equation:
x + x+10 + 2x-30 = 180
4x = 200
-------
x = 50 degrees
x+10 = 60 degrees
2x-30 = 70 degrees
===============
Cheers,
Stan H.
==============